# Hierarchical Learning-Enhanced MPC for Safe Crowd Navigation with Heterogeneous Constraints

### v1

@@ III-A1 Problem formulation
Consider a episodic environment in a 2D manifold $\mathcal{W}\in\mathbb{R}^{2}$, where an autonomous mobile robot must navigate from an initial state $\mathbf{x}_{0}$ to a goal $\mathbf{g}$. During this process, the robot needs to interact with pedestrians and heterogeneous static obstacles. Motion planning of the robot can be formulated as an optimal control problem:
@@ III-C Neural Network Architecture
Building upon the classical actor-critic architecture, we design the policy and value networks using a graph neural network layer as a feature extractor. Based on empirical results, we do not employ a shared feature extraction layer as similar works do. The output of each GNN is concatenated with the robot’s node feature to form the integrated fixed-length observation $\mathbf{H}$. Both the policy network $\pi_{\theta}$ and the value network $V$ consist of two hidden fully connected layers (FCLs) with 256 rectified linear units.
@@ IV-B2 Ablation study
To validate this proposition, we designed an additional set of ablation experiments. The entire frontend network was removed, and the A* algorithm was utilized to generate a reference trajectory, considering only structured obstacles. Subsequently, the same MPC-based backend was employed for trajectory refinement, and the local goal was selected from the reference trajectory based on the length of the MPC prediction horizon and the robot’s reference velocity. The experimental results, as summarized in Table IV, clearly show that without the guidance of the frontend network, the backend must handle complex interactions more frequently. The S-T joint search module significantly enhances navigation performance by ?, demonstrating its effectiveness.
Furthermore, we investigated the impact of the discretization resolution of the robot’s neighborhood on overall performance, with the results presented in Figure ?. As the discretization resolution increases, the planner’s performance improves correspondingly and stabilizes after exceeding 81 discrete points. Based on this observation, we set the action space dimension of the policy network to $D=81$ for all other experiments.
@@ IV-C Qualitative Evaluation
To provide a more intuitive illustration of the proposed method’s characteristics, we selected several representative simulation scenarios to visualize the navigation process, as shown in Figure ?.
@@ V-B Comparative Results
With a relatively high degree of repeatability, we designed a corridor-like experimental scenario consisting of one large polygonal obstacle, five static circular obstacles, and three pedestrians. The distance between the initial and target positions was set to 8 m, consistent with the simulation environment. The proposed method was evaluated against SAGE and HGAT-DRL, with each method tested 20 times. Statistical results are summarized in Table V. It is important to note that, due to the inherent non-reproducibility of pedestrian behaviors, these experimental results, while providing an intuitive reference, should not be considered strictly rigorous.
@@ V-C Quantitative Results
To provide an intuitive illustration of the experimental results, several representative scenarios from the experiments were selected for visualization, as shown in Fig. 4.
Figure 4: Temporal visualizations of typical experimental scenarios, including synchronized perception and planning results. More experimental results can be found in the video attachment.
@@ VI Conclusion and Discussion
Figure 5: Visualized results of long-distance multi-point navigation experiments. The lettered labels in the figure indicate representative frames.

### v2

@@ III-A1 Problem formulation
Consider an episodic environment in a 2D manifold $\mathcal{W}\in\mathbb{R}^{2}$, where an autonomous mobile robot must navigate from an initial state $\mathbf{x}_{0}$ to a goal $\mathbf{g}$. During this process, the robot needs to interact with pedestrians and heterogeneous static obstacles. Motion planning of the robot can be formulated as an optimal control problem:
@@ III-C Neural Network Architecture
Building upon the classical actor-critic architecture, we design the policy and value networks using graph neural network layers as the feature extractors, as illustrated in Fig. 2. Based on empirical results, we do not employ a shared feature extraction layer as similar works do. The output of each GNN is concatenated with the robot’s node feature to form the integrated fixed-length observation $\mathbf{H}$. Both the policy network $\pi_{\theta}$ and the value network $V$ consist of two hidden fully connected layers (FCLs) with 256 rectified linear units.
Figure 2: An illustration of the architecture of the proposed GNN-based policy network with invalid action masking mechanism.
@@ IV-B2 Ablation study
To validate this proposition, we designed an additional set of ablation experiments. The entire frontend network was removed, and the A* algorithm was utilized to generate a reference trajectory, considering only structured obstacles. Subsequently, the same MPC-based backend was employed for trajectory refinement, and the local goal was selected from the reference trajectory based on the length of the MPC prediction horizon and the robot’s reference velocity. The experimental results, as summarized in Table IV, clearly show that without the guidance of the frontend network, the backend must handle complex interactions more frequently. The S-T joint search module significantly enhances the navigation success rate by 8% without sacrificing efficiency, demonstrating its effectiveness.
Furthermore, we investigated the impact of the discretization resolution of the robot’s neighborhood on overall performance, with the results presented in Table III. As the discretization resolution increases, the planner’s performance improves correspondingly and stabilizes after exceeding 81 discrete points. Based on this observation, we set the action space dimension of the policy network to $D=81$ for all other experiments.
@@ IV-C Qualitative Evaluation
To intuitively demonstrate the characteristics of the proposed method, several representative simulation scenarios are selected to visualize the navigation processes of HALO and two baseline methods, SAGE and HGAT-DRL, as shown in Fig. 5. A qualitative analysis of the illustrated navigation processes reveals that, while all three methods exhibit a certain level of scene understanding and can select less interactive routes based on pedestrian motion trends, only the proposed method is capable of correctly guiding the robot to bypass obstacles from the right and successfully reach the goal with a smooth and highly consistent trajectory in the complex scenarios depicted.
Figure 5: Qualitative comparison of the navigation process in representative simulation scenario. The ego-robot is depicted in black hollow circle with a red arrow indicates its heading direction and a red asterisk marks its goal. Pedestrians are depicted in blue hollow circles, static obstacles are depicted in gray solid circles and structured obstacles such as walls are plotted in black thin lines. Specifically for HALO, blue dots represent the local goals generated by the frontend network, red dots denote the invalid local goals masked out by the action mask, and red thin lines correspond to the locally optimal trajectories generated by the backend module.
@@ V-B Comparative Results
With a relatively high degree of repeatability, we designed a corridor-like experimental scenario consisting of one large polygonal obstacle, five static circular obstacles, and three pedestrians. The distance between the initial and target positions was set to 8 m, consistent with the simulation environment. The proposed method was evaluated against SAGE and HGAT-DRL, with each method tested 20 times. Statistical results are summarized in Table V. To better highlight the planner’s capabilities, the experimental scenarios are intentionally designed to be more challenging than typical daily-life environments. Furthermore, the participating pedestrians are instructed to avoid actively yielding to the robot whenever possible. It is important to note that, due to the inherent non-reproducibility of pedestrian behaviors, these experimental results, while providing an intuitive reference, should not be considered strictly rigorous.
@@ V-C Quantitative Results
To provide an intuitive illustration of the experimental results, several representative scenarios from the experiments were selected for visualization, as shown in Fig. 6. First, a simple corridor scenario is used to illustrate the planner’s anticipatory decision-making capability in Fig. 6(a). Under identical initial and goal conditions, the planner is able to select different topological paths to bypass the obstacle from either side, based on the pedestrians’ motion trends, thereby reducing interactions. In Fig. 6(b), a slow-moving pedestrian is walking in the same direction ahead of the robot, blocking its path. Additional obstacles further narrow the traversable area, limiting the robot’s options for detour. In this scenario, the robot is able to follow the slow pedestrian at a similar speed and resume normal speed to reach the goal once the path is cleared.
Figure 6: Temporal visualizations of typical experimental scenarios, including synchronized perception and planning results. Thin lines with colors corresponding to the dynamic obstacles indicate their explicitly predicted future trajectories. More experimental results can be found in the video attachment.
Fig. 6(c) presents five sequential snapshots from a typical episode in the comparative experiments. Initially, with pedestrians present on both sides, the robot performs a backward motion to avoid interfering with the blue-marked pedestrian moving leftward and is guided to bypass the rectangular obstacle from the right side. Subsequently, when the right passage becomes occupied by a green-marked pedestrian, the robot proactively backs off to avoid collision, then handles interactions with two upcoming pedestrians by waiting. In the final snapshot, the robot performs another backward maneuver to avoid a pedestrian deliberately moving toward it and safely reaches the goal.
In Fig. 6(d), three sequential snapshots depict another scenario. The planner first guides the robot to pass on the right to avoid interaction with the blue-marked pedestrian, based on the pedestrian’s motion trend. As the robot approaches a bench, its feasible path is blocked by a red-marked pedestrian with no possible detour. The robot slows down to wait until the pedestrian passes and sits down, then detours to the left to reach the goal. A similar deceleration is executed in the final snapshot, ensuring safe arrival at the destination.